Morning:
  • JZ fixes a fuse that was burnt because of the new piezo electronics required too much power.
  • Checking for the poweron - poweroff procedures from arbitrator if a fuse fails the system could not be switched down...
  • Bug fixes in the AARB (ask JB for details)

13:00: switchon LgswSystem SX and DX, we try to revive the APDs...

Both SX and DX systems fail to close the loop, the TT diverges.

We focus on SX: the APD3 fiber seems misconnected or broken, we go up to check:

APD Ch Rack AGW
1 0 F4 F4
2 1 F3 F3
3 2 Spare F1
4 3 F2 F2

We try to acquire a new dark, but now also ADP2 seems to not work properly: with flat illumination the counts are a factor 100 less than other channels.
Topic revision: r3 - 01 May 2017, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback